OpenStudy (anonymous):

Let f(x) = 4x - 5 and g(x) = 6x - 3. Find F(x) - g(x). A. 2x - 5 B. 2x + 5 C. -2x - 2 D. -2x - 8

OpenStudy (lncognlto):

f(x) - g(x) simply equals the function f minus the function g. therefore, f(x) - g(x) = (4x - 5) - (6x - 3) = 4x - 5 - 6x + 3 = ?

OpenStudy (anonymous):

is it c? @lncognlto

OpenStudy (lncognlto):

Yep, c is correct. ^-^